Find the quantity of heat required to melt completely 20g of lead at 27 degree
1
Expert's answer
2020-09-29T09:40:48-0400

We need first to heat up lead to melting point temperature ("327.5^\circ C") and then some energy needed for phase transition. So total quantity of heat will be

"Q = cm \Delta T+ \lambda m = 140 \cdot 20 \cdot 10^{-3} (327.5 -27) + 0.25 \cdot 10^5 \cdot 20 \cdot 10^{-3} = 841.4 + 500 = 1341.4 \; J"

where the first term is heating to melting point, the second term - energy needed for melting; c is specific heat capacity, "\lambda" is heat of fusion.

Answer: "1341.4 \; J"